### v1

@@ Abstract
Methods. Following Dainotti et al. (2024c), we have taken a second step to overcome this issue by adding 30 more GRBs to our ensemble supervised machine learning training sample, an increase of 20%, which will help us obtain better redshift estimates. In addition, we have built a freely accessible and user-friendly web app that infers the redshift of long GRBs (LGRBs) with plateau emission using our machine learning model. The web app is the first of its kind for such a study and will allow the community to obtain redshift estimates by entering the GRB parameters in the app.
@@
Conclusions. The results presented in this research will enable the community to increase the sample of GRBs with known redshift estimates. This can help address many outstanding issues, such as GRB formation rate, luminosity function, and the true nature of low-luminosity GRBs, and enable the application of GRBs as standard candles.
@@ 1 Introduction
To address the challenges outlined above, we propose here a machine learning (ML) based approach for estimating the redshift of GRBs. This work is a follow-up of Dainotti et al. (2024c) and Dainotti et al. (2024b) and utilizes an enlarged sample, 20% larger than the sample in Dainotti et al. (2024c), to train our ML models. In addition, we have extended the generalization sample by 35%. This is the set of GRBs without known redshifts, and we predict the redshift of 276 GRBs using our trained ML model. Furthermore, we release a user-friendly web application that provides redshift estimates for new GRBs, built on the ML models described in this paper.
@@ 2 Advantage of machine learning VS linear relationships
The extension of the Dainotti relations in three dimensions has also been achieved by Xu & Huang (2012), where they add the isotropic $\gamma$-ray energy to the two-dimensional relation, naming it the $L-T-E$ correlation. Zhao et al. (2019); Tang et al. (2019) independently confirmed the existence of this correlation for 174 GRBs from Swift. Tang et al. (2019) also improved the best-fit relation to be $L_{a}\propto T_{a}^{-1.01}E^{0.84}$. This work was expanded upon further by Xu et al. (2021) and Deng et al. (2023). Xu et al. (2021) showed the redshift evolution of the $L-T-E$ correlation and introduced a new correlation with the spectral peak energy labeled $L-T-E_{p}$ correlation. Deng et al. (2023) updated coefficients of the $L-T-E$ correlation of Tang et al. (2019) to $L_{a}\propto T_{a}^{-0.99}E^{0.86}$ with 210 LGRBs with plateau and also derived pseudo-redshifts for 108 plateau GRBs. However, their redshift estimates are limited by intrinsic scatter and functional form of the $L-T-E$ correlation. Moreover, the precision of their approach is not tested against GRBs with known redshift. In comparison, our ML-based approach tests the performance of our model on events with known redshifts.
@@
Another issue, which is often overlooked, is that these pseudo-redshifts are inherently subject to inaccuracies. This is because of their non-linear relation to the luminosity distance ($D_{L}$).
The major problem in the forward-fitting relationships is the dependence on cosmology. Indeed, one of the parameters of this relationship depends on the $D_{L}$, which itself depends on cosmological parameters. Thus, using any of these methods would induce a circularity problem, which must be avoided, especially when we tackle cosmology. In addition, given the functional form of $\log(D_{L})$, as it flattens at $z\gtrsim 3$, any small variation of $\log(D_{L})$ can become a large variation in redshift determination. Hence, the difficulty in determining the redshifts accurately with these forward fitting methods. Furthermore, based on relationships, the forward fitting methods mentioned do not account for the errors on the parameters, thus underestimating the realistic uncertainties on the redshifts. Thus, we have followed a route via ML algorithms to obtain more accurate predictions and overcome these issues.
@@ 3 The Data Sample
3.1 Comparison between the previous sample and the extended sample
In Fig.2 we present the scatter matrix plot between the training set from Dainotti et al. (2024c) (cyan) and the 30 new GRBs (orange) added in this analysis for comparing the distribution of the variables. We applied the Kolmogorov-Smirov test (KS-test) for each variable between the two data sets. All the variables pass the KS-test with a p-value $>$ 0.05, indicating that they originate from the same parent population. However, $T_{90}$ and $\gamma$ obtained p-values $<$ 0.05.
In the left panel of Fig. 3, the redshift distribution is illustrated, with the new sample depicted by a dashed line and the old sample by a solid line, both normalized to 1. The right panel of Fig. 3 presents the actual distribution. The redshift distributions pass the KS-test with a p-value of 0.35. Thus, we can say that the two sets of redshifts originate from the same parent population.
@@ 4 Methodology and web app
In this research, we expand the methodology implemented in Dainotti et al. (2024c), and in addition, we make it available to the community via an easy-to-use and open-source web app. This web app is designed to let users input GRB parameters and quickly obtain redshift estimates within hours of the GRB trigger. This will allow the community to conduct follow-up observations of interesting bursts, such as high-$z$ events or low-$z$ low-luminosity GRB (LLGRB) events. This web app is also highly modular, and users are free to train their own custom models using their data.
We here detail in each subsection the various methods implemented in our analysis and the associated functions of the web-app.
@@ 4.1 MICE
To minimize data loss and enhance the size of both the training and generalization sets, we utilized the imputation technique known as Multivariate Imputation by Chained Equations (MICE, Van Buuren & Groothuis-Oudshoorn (2011); Gibson et al. (2022)). MICE leverages the data currently in use to infer the missing data. The method applied within MICE is called Predictive Mean Matching (PMM, Little & Rubin (2019)). PMM initially fills the missing entries with the mean value of that particular feature. Here, we stress that random imputation is appropriate since the missing values of the variable are mainly due to the orbital gaps in the light curves. Then, it trains a regression model on the available data and, using this model, predicts the missing entries. Next, it finds the feature’s closest observed values to this prediction and randomly selects one of them as the imputed value. This process is carried out iteratively for each feature with missing entries until the values stabilize or a predetermined number of iterations is reached. We perform this iteration 20 times to have a reliable imputation. With the application of MICE, we increase our training sample by 14% and the generalization set by 4%.
@@ 4.2.2 The ML Models used
We tested multiple different ML models to find the best set of models that can give accurate redshift predictions. Based on our analysis, we determined three models that perform best, which are described below.
@@ 4.4 The Superlearner
SuperLearner is an ensemble method that combines multiple ML models and leverages these models to obtain the best prediction. SuperLearner assigns to each model a weight based on its performance, and the sum of all weights is 1. The performance of the constituent models is determined via the 10fCV method, which is built into SuperLearner. The advantage of the SuperLearner is that its prediction is a weighted average of the predictions of its constituent models. We also use SuperLearner weights to select models that perform the best and models that obtain higher weights are included in the final nested 100 times 10fCV procedure. This procedure provides us with the results presented in this paper.
@@ 4.6 Bias correction
In this technique, a linear model is fitted to the sorted values of observed and predicted $\log(z+1)$. The predicted values are corrected by multiplying the slope and adding the intercept of the linear fit. We perform this bias correction on the 10fCV results, and the generalization set predictions.
@@ 5.3 Estimating new redshifts in the generalization sample
It should also be noted that GRBs with and without $z_{obs}$ might differ in their physical origin and chances of observations. Thus, they can have different intrinsic redshift distributions. The low-luminosity sources have a lower chance of having their redshift measured due to the difficulty of obtaining the redshift from the afterglow spectral features, the trouble of observing the host galaxy, or the fact that the host galaxy is not in the database and is not pointed by any observing facilities. Moreover, we observe a statistically smaller $\log(NH)$ in the generalization set. This corresponds to less absorption via the intergalactic medium, thus making the absorption-based redshift measurements harder. Those two effects could explain the difference in redshift distributions between the training and generalization sets.
We show a boxplot distribution of the predicted redshifts in the lower panel of Fig. 5.3. The red and grey boxplots show the bias-corrected and non-bias-corrected distributions. The x-axis shows the names of the generalization set GRBs, and the y-axis shows the redshift values. Note that because of the image’s scale, not all the GRB names are present along the axis. Our work here has enabled us to increase the Swift sample with redshift estimates by 110%.
@@ 5.4.1 Future of Multi-wavelength Redshift Estimates via Machine Learning
For the optical afterglow,e have gathered 536 LCs through the 416 ground-based Telescopes, the HST, the Swift/UVOT satellites, the GCN, and private communications. These 536 GRBs, observed from 05/1997 to 10/2023, have $z_{obs}$ and optical afterglow. Out of these 536 GRBs, 180 (34%) GRBs have optical plateau observations. As of this writing, according to the Greiner webpage [^3] https://www.mpe.mpg.de/ jcg/grbgen.html, the total number of GRBs with optical afterglow is 932, and 622 have secure redshifts. Thus, 310 GRBs (932 - 622) do not have redshift but have an optical afterglow. Assuming the same fraction of plateau observations in the optical afterglow, our current generalization set would be 34% of 310, which is 104 GRBs. If we extrapolate to 09/2025, we expect to have 112 GRBs with an optical plateau and no measured redshifts. Similarly, we expect to have 193 GRBs by 09/2025, which have an optical plateau and $z_{obs}$. Thus, we expect $>$112 GRBs with $z_{pred}$ from ML analysis. Thus, the new $z_{pred}$ will be $>$436 (324+112), where 324 and 112 are GRBs with X-ray and optical plateaus, respectively. With this increase of 92% in the redshift sample, we will enhance the accuracy of the LF and density rate evolution estimates, thus helping to address the discrepancy between the GRBFR and SFR evolution. The above numbers are also reflected in the Table 1. This analysis can be extended for future reference, and it is our basis to compute how many GRBs we will have approximately in the next one, two, and three years from now.
@@
According to Table.9 of Dainotti et al. (2022), to reach the precision of Conley et al. (2011), the lowest number of X-ray platinum sample GRBs with redshift needed, with half error bars on plateau parameters, is 354. To achieve this number, we assume that by using an updated lightcurve reconstruction (LCR) technique, an improvement of the methodology from Dainotti et al. (2023), which is in preparation, we will have the same number of GRBs with half error bars (denoted by n=2) as with full error bars (denoted by n=1). Currently, we have 105 GRBs that belong to the platinum sample with $z_{obs}$. Assuming the same fraction of platinum GRBs exists in the generalization set, we should have another 125 GRBs from the generalization set used in this work. Combining these two, we estimate that we currently have 230 platinum sample GRBs with redshift estimates. Further, considering the observations we will obtain from SVOM (Godet et al. (2014) and SVOM webpage [^4] https://www.eoportal.org/satellite-missions/svom-eclairs-gamma-ray-imager) and the application of ML and LCR, we can expect to reach the precision of Conley et al. (2011) by 2029 (see PLATtrim20 column of Table.2). Thus, this ML effort reduces the number of years required to reach the precision of Conley et al. (2011) by 26 years. Similarly, to reach the precision of Betoule et al. (2014), we need the observational capabilities of THESEUS. According to our estimates in Table.2, again, using LCR and ML to boost the numbers and GRBs with half error, we should be able to reach the precision of Betoule et al. (2014) by 2041 at least.

### v2

@@ Abstract
Methods. Following Dainotti et al. (2024c), we have taken a second step to overcome this issue by adding 30 more GRBs to our ensemble supervised machine learning training sample, an increase of 20%, which will help us obtain more accurate pseudo-redshifts. In addition, we have built a freely accessible and user-friendly web app that infers the redshift of long GRBs (LGRBs) with plateau emission using our machine learning model. The web app is the first of its kind for such a study and will allow the community to obtain pseudo-redshifts by entering the GRB parameters in the app.
@@
Conclusions. The results presented in this research will enable the community to increase the sample of GRBs with known pseudo-redshifts. This can help address many outstanding issues, such as GRB formation rate, luminosity function, and the true nature of low-luminosity GRBs, and enable the application of GRBs as standard candles.
@@ 1 Introduction
To address the challenges outlined above, we propose here a machine learning (ML) based approach for estimating the redshift of GRBs. This work is a follow-up of Dainotti et al. (2024c) and Dainotti et al. (2024b) and utilizes an enlarged sample, 20% larger than the sample in Dainotti et al. (2024c), to train our ML models. In addition, we have extended the generalization sample by 35%. This is the set of GRBs without known redshifts, and we predict the redshift of 276 GRBs using our trained ML model. Furthermore, we release a user-friendly web application that provides pseudo-redshifts for new GRBs, built on the ML models described in this paper.
@@ 2 Advantage of machine learning VS linear relationships
The extension of the Dainotti relations in three dimensions has also been achieved by Xu & Huang (2012), where they add the isotropic $\gamma$-ray energy to the two-dimensional relation, naming it the $L-T-E$ correlation. Zhao et al. (2019); Tang et al. (2019) independently confirmed the existence of this correlation for 174 GRBs from Swift. Tang et al. (2019) also improved the best-fit relation to be $L_{a}\propto T_{a}^{-1.01}E^{0.84}$. This work was expanded upon further by Xu et al. (2021) and Deng et al. (2023). Xu et al. (2021) showed the redshift evolution of the $L-T-E$ correlation and introduced a new correlation with the spectral peak energy labeled $L-T-E_{p}$ correlation. Deng et al. (2023) updated coefficients of the $L-T-E$ correlation of Tang et al. (2019) to $L_{a}\propto T_{a}^{-0.99}E^{0.86}$ with 210 LGRBs with plateau and also derived pseudo-redshifts for 108 plateau GRBs. However, their pseudo-redshifts are limited by intrinsic scatter and functional form of the $L-T-E$ correlation. Moreover, the precision of their approach is not tested against GRBs with known redshift. In comparison, our ML-based approach tests the performance of our model on events with known redshifts.
@@
Another often overlooked issue is the inherent inaccuracies of these pseudo-redshifts, arising from their non-linear dependence on the luminosity distance ($D_{L}$).
[…]
The major problem in the forward-fitting relationships is the dependence on cosmology. Indeed, one of the parameters of this relationship depends on the $D_{L}$, which itself depends on cosmological parameters. We also would like to stress that it is important to have a method that is not dependent on a given cosmological model; otherwise, we incur the so-called ”circularity problem”. Indeed, it would be highly circular to use the pseudo-redshifts derived from the forward fitting method, which assumes a given cosmology to then derive cosmological parameters. Even if the goal of the pseudo-redshift is not to derive the cosmological parameters, we avoid being biased toward a particular or a given cosmological model for the analysis of the luminosity function and density rate evolution here. Although, the errorbars for the luminosity function are not negligible, in the procedure of the selection biases detailed in (Dainotti et al., 2024b) if the cosmology had been inferred, we would have had a different number of associated sets (described in the appendix of that paper), because of the choice of a different limiting luminosity. This, hence, will also affect the density rate evolution.
Thus, using any of these methods would induce a circularity problem, which must be avoided, especially when we tackle cosmology. In addition, given the functional form of $\log(D_{L})$, as it flattens at $z\gtrsim 3$, any small variation of $\log(D_{L})$ can become a large variation in redshift determination. Hence, the difficulty in determining the redshifts accurately with these forward-fitting methods. Furthermore, based on relationships, the forward fitting methods mentioned do not account for the errors on the parameters, thus underestimating the realistic uncertainties on the redshifts. Thus, we have followed a route via ML algorithms to obtain more accurate predictions and overcome these issues.
@@ 3 The Data Sample
In Fig.2 we present the scatter matrix plot between the training set from Dainotti et al. (2024c) (cyan) and the 30 new GRBs (orange) added in this analysis for comparing the distribution of the variables. We applied the Kolmogorov-Smirov test (KS-test) and the Anderson-Darling test for each variable between the two data sets.
All the variables pass the two tests with a p-value $>$ 0.05, indicating that they originate from the same parent population, except $T_{90}$ and $\gamma$, which obtain p-values $<$ 0.05 for both KS-test and AD-test, and $\log(F_{a})$ which obtained a p-value $<$ 0.05 for AD-test. For further discussion see the Appendix and the Fig. 19. Here, we are presenting the comparison between the redshift distributions.
In the left panel of Fig. 3, the redshift distribution is illustrated, with the new sample depicted by a dashed line and the old sample by a solid line, both normalized to 1. The right panel of Fig. 3 presents the actual distribution with no normalization. The redshift distributions pass the KS-test with a p-value of 0.35 and the AD-test with p-value of 0.41, as it can be seen from the right panel of Fig. 3 . Thus, we cannot reject the hypothesis that the two sets of redshifts originate from the same parent population.
@@ 4 Methodology and web app
In this research, we expand the methodology implemented in Dainotti et al. (2024c), and in addition, we make it available to the community via an easy-to-use and open-source web app. This web app is designed to let users input GRB parameters and quickly obtain pseudo-redshifts within hours of the GRB trigger. This will allow the community to conduct follow-up observations of interesting bursts, such as high-$z$ events or low-$z$ low-luminosity GRB (LLGRB) events. This web app is also highly modular, and users are free to train their own custom models using their data.
We here detail in each subsection the various methods implemented in our analysis and the associated functions of the web-app. The webapp is available for free on the following Github page:https://github.com/gammarayapp/GRB-Web-App
@@ 4.1 MICE
To minimize data loss and enhance the size of both the training and generalization sets, we utilized the imputation technique known as Multivariate Imputation by Chained Equations (MICE, Van Buuren & Groothuis-Oudshoorn (2011); Gibson et al. (2022)). MICE leverages the data currently in use to infer the missing data. Here, for the data imputation to be reliable, MICE implicitly assumes that the data is missing at random. This means that the observed data can explain the missing data, and the probability of the data not being observed does not depend on the missing data’s value itself. Furthermore, MICE also assumes that the data have a multivariate distribution and, thus, considers the relations between features for imputing missing data. Here, we stress that random imputation is appropriate, as the missing values of the variable are mainly due to the good and bad time intervals, orbital gaps, instrumental errors, and satellite follow-up repointing.
The imputed data are sampled from the observed data, and the distribution is kept the same. We did this on purpose so that we would not have differences that could bias the analysis. In this study, we used the predictive mean matching method (PMM), specifically the mida-touch approach (see Little & Rubin (2019) for details). Using PMM ensures that the imputed data is sampled from the observed data. This method fills in a feature’s missing values with its mean and then iteratively refines these values by training a model on the complete data available. Each prediction is assigned a probability based on its proximity to the imputed value for the missing variable. The missing entry is then set by randomly choosing it from the observed values of the respective variable, weighted according to the previously defined probability. This process is carried out iteratively for each feature with missing entries until the values stabilize or a predetermined number of iterations is reached. We perform this iteration 20 times to have a reliable imputation. With the application of MICE, we increase our training sample by 14% and the generalization set by 4%.
@@ 4.2.2 The ML Models used
ML models in general require a large amount of data to train. Here, even though we have the largest sample of LGRBs with plateau and redshift, this is still at least an order of magnitude less than the amount of data required to reliably train more complex ML models, such as neural networks. Thus, we tested multiple different ML models to find the best set of models that can give accurate redshift predictions with the data set currently available. The three chosen models performed the best during our analysis. Furthermore, to test whether the performance of these models is maintained as the dataset increases, we have performed a Monte Carlo simulation as well (see Sec.4.7). Thus, based on our analysis we are confident that these chosen models are best suited for the data at hand. .
@@ 4.2.3 Formula Generation
Our approach of using analytic expressions is similar to the symbolic regression approach of Cranmer (2023). However, the key difference here is that we are not using exponential powers or trigonometric functions of our features. Given our smaller data sample, we focus on additive expressions. However, investigating a symbolic regression could provide an interesting avenue of research. For now, it is beyond the scope of this paper.
@@ 4.4 The Superlearner
SuperLearner is an ensemble method that combines multiple ML models and leverages these models to obtain the best prediction. SuperLearner assigns to each model a weight based on its performance (larger weight for better performance), and the sum of all weights is 1, as indicated in the following formula:
$\sum{W_{i}}=1$
where $W_{i}$ is the weight for the $i^{th}$ model. The performance of the constituent models is determined via the 10fCV method, which is built into SuperLearner. The advantage of the SuperLearner is that its prediction is a weighted average of the predictions of its constituent models, with better-performing models obtaining larger weights. However, different ML models have their own individual implicit assumptions and priors, which can impact the redshift estimation. Here, we use the SuperLearner weights to select models whose implicit priors best work with our dataset. Models that obtain higher weights are included in the final nested 100 times 10fCV procedure, which provides the results presented in this paper. Furthermore, we are combining parametric (GLM), semi-parametric (GAM) and non-parametric (RF) into the final SuperLearner model, highlighting the benefits of an ensemble approach. A linear model like GLM enables higher performance for linear relations in the data, while GAM and RF enable better fit for non-linear relations. Thus, the models share features along with their individual features, enabling a performative ensemble ML model.
@@ 4.6 Bias correction
In this technique, first, the observed and predicted values of the response variable are sorted in ascending order. Next, a linear model is fitted between them and the predicted values are corrected by multiplying the slope ($A)$ and adding the intercept ($B$) of the linear fit. Furthermore, bias correction is performed separately in three redshift ranges: $20<z<$ 2, $3.52<z<$ 3.5, and $8.43.5<z<$ 8.4. These particular ranges were selected as they produced the best results for our particular ML model and dataset. The bias correction equation is as follows:
$Z_{C}=B+A*Z_{P}$
where $Z_{C}$ is the corrected $\log(z+1)$ prediction, $Z_{P}$ is the SuperLearner prediction. The assumption we are making here is that the $\log(z_{pred}+1)$ follows the same distribution as $\log(z_{spec}+1)$.
@@ 5.3 Estimating new redshifts in the generalization sample
It should also be noted that GRBs with and without $z_{obs}$ might differ in their physical origin and chances of observations. Thus, they can have different intrinsic redshift distributions. The low-luminosity sources have a lower chance of having their redshift measured due to the difficulty of obtaining the redshift from the afterglow spectral features, the trouble of observing the host galaxy, or the fact that the host galaxy is not in the database and is not pointed by any observing facilities. Moreover, we observe a statistically smaller $\log(NH)$ in the generalization set. This corresponds to less absorption via the intergalactic medium, thus making the absorption-based redshift measurements harder. Those two effects could explain the difference in redshift distributions between the training and generalization sets. We discuss the physical origins of each variable in Dainotti et al. (2025). However, we can adopt certain strategies to help us constrain the differences. The first strategy one could adopt would be to ensure that the parameter spaces of the generalization set and training set are same, which is what we do here in this paper. Indeed, to prevent biases in the pseudo-redshifts, we trim the generalization set before predicting its redshift so that our ML model will predict only those GRBs whose parameter space falls within our training set. This will ensure that any outlier GRBs in the generalization set that are more likely to have a different physical origin can be minimized. Another strategy would be better classify GRBs, regardless of whether they have $z$ or not. Indeed, some of us have already attempted to address this Bhardwaj et al. (2023), where we attempted to classify GRBs based on their properties using unsupervised ML. Efforts in this direction can help ensure that a homogeneity is maintained between the training set of an ML model and the generalization set used. Next, we have already investigated how the performance of our ML model changes when more GRBs are added to the generalization set, making sure that our model is predictive enough to handle the additional data and continue to perform at the expected level. We explore this in the next section.
We show a boxplot distribution of the predicted redshifts in the lower panel of Fig. 5.3. The red and grey boxplots show the bias-corrected and non-bias-corrected distributions. The x-axis shows the names of the generalization set GRBs, and the y-axis shows the redshift values. Note that because of the image’s scale, not all the GRB names are present along the axis. Our work here has enabled us to increase the Swift sample with pseudo-redshifts by 110%.
@@ 5.4.1 Future of Multi-wavelength Redshift Estimates via Machine Learning
For the optical afterglow, we have gathered 536 LCs through the 416 ground-based Telescopes, the HST, the Swift/UVOT satellites, the GCN, and private communications. These 536 GRBs, observed from 05/1997 to 10/2023, have $z_{obs}$ and optical afterglow. Out of these 536 GRBs, 180 (34%) GRBs have optical plateau observations. As of this writing, according to the Greiner webpage [^3] https://www.mpe.mpg.de/ jcg/grbgen.html, the total number of GRBs with optical afterglow is 932, and 622 have secure redshifts. Thus, 310 GRBs (932 - 622) do not have redshift but have an optical afterglow. Assuming the same fraction of plateau observations in the optical afterglow, our current generalization set would be 34% of 310, which is 104 GRBs. If we extrapolate to 09/2025, we expect to have 112 GRBs with an optical plateau and no measured redshifts. Similarly, we expect to have 193 GRBs by 09/2025, which have an optical plateau and $z_{obs}$. Thus, we expect $>$112 GRBs with $z_{pred}$ from ML analysis. Thus, the new $z_{pred}$ will be $>$436 (324+112), where 324 and 112 are GRBs with X-ray and optical plateaus, respectively. With this increase of 92% in the redshift sample, we will enhance the accuracy of the LF and density rate evolution estimates, thus helping to address the discrepancy between the GRBFR and SFR evolution. The above numbers are also reflected in the Table 1. This analysis can be extended for future reference, and it is our basis to compute how many GRBs we will have approximately in the next one, two, and three years from now.
@@
According to Table.9 of Dainotti et al. (2022), to reach the precision of Conley et al. (2011), the lowest number of X-ray platinum sample GRBs with redshift needed, with half error bars on plateau parameters, is 354. To achieve this number, we assume that by using an updated lightcurve reconstruction (LCR) technique, an improvement of the methodology from Dainotti et al. (2023), which is in preparation, we will have the same number of GRBs with half error bars (denoted by n=2) as with full error bars (denoted by n=1). Currently, we have 105 GRBs that belong to the platinum sample with $z_{obs}$. Assuming the same fraction of platinum GRBs exists in the generalization set, we should have another 125 GRBs from the generalization set used in this work. Combining these two, we estimate that we currently have 230 platinum sample GRBs with pseudo-redshifts. Further, considering the observations we will obtain from SVOM (Godet et al. (2014) and SVOM webpage [^4] https://www.eoportal.org/satellite-missions/svom-eclairs-gamma-ray-imager) and the application of ML and LCR, we can expect to reach the precision of Conley et al. (2011) by 2029 (see PLATtrim20 column of Table.2). Thus, this ML effort reduces the number of years required to reach the precision of Conley et al. (2011) by 26 years. Similarly, to reach the precision of Betoule et al. (2014), we need the observational capabilities of THESEUS. According to our estimates in Table.2, again, using LCR and ML to boost the numbers and GRBs with half error, we should be able to reach the precision of Betoule et al. (2014) by 2041 at least.